Sample Problems for Basic Dosage Calculation

Order: Amoxicillin 0.25 g p.o. every 8 hours. Available: Amoxicillin 125mg tablets.

How many tablets will the nurse give per dose?

Order: Zofran 8 mg p.o.t.i.d. Available: Zofran in a 100mL bottle labeled 4 mg/tsp.

How many mL will the nurse administer for each dose?

Give Fortaz 50 mg/kg p.o. t.i.d. to a child who weighs 25.5 kg. Fortaz is available in an oral

suspension labeled 100mg/mL. How many mL would the nurse administer per dose?

Give Ceclor 45 mg/kg/day p.o. in 3 divided dosesfor a patient who weighs 66 pounds. A 75 mL

stock medication is labeled Ceclor 125mg/mL. How many mL would the nurse administer per

dose?

The doctor ordered Magnesium Sulfate 4gm IV to be given over 20 minutes. Magnesium Sulfate come in 40gm/1000mL IV bottle. You would set your pump at _____mL/hour.

A. 250 mL/hr

B. 300 mL/hr

C. 350 mL/hr

D. 380 mL/hr

The doctor ordered Magnesium Sulfate 2gm/ hour IV. Magnesium Sulfate come in 40gm/1000mL IV bottle. You would set your pump at _____mL/hour.

A. 50 mL/hr

B. 60 mL/hr

C. 55 mL/hr

D. 53 mL/hr

The physician ordered Ritodrine IV 70mcg/min. The pharmacy sent up Ritodrine 150mg premixed in 500mL D5W. You would set your IV pump at ________mL/hour.

A. 15 mL/hr

B. 16 mL/hr

C. 13 mL/hr

D. 14 mL/hr

The physician ordered Ritodrine IV 50 mcg/min. The pharmacy sent up Ritodrine 150 mg premixed in 500mL D5W. You would set your IV pump at ________mL/hour.

A. 9 mL/hr

B. 10 mL/hr

C. 11 mL/hr

D. 12 mL/hr

The physician ordered the Pitocin infusion to run at 6 mu/min. The pharmacy sent up 10 units of Pitocin in 500 mL of D5LR. You would set your pump at _____________ mL/hour.

A. 18 mL/hr

B. 19 mL/hr

C. 20 mL/hr

D. 21 mL/hr

The medications scheduled for your patient include Keflex 1.5 grams in 50 ml of a 5% Dextrose solution. According to the pharmacy, this preparation should be administered in 30 minutes. You would set your IV pump at __________mL/hour.

A. 100 mL/hr

B. 110 mL/hr

C. 115 mL/hr

D. 120 mL/hr
Answer #1

0.25 g = 250 mg

Number of tablets = 250/125 = 2

4 mg in one tsp (5 ml)

8 mg in 10 ml.

So nurse should give 10 ml per dose.

Dose = 50 * 25.5 = 1275 mg

100 mg per ml

1275 mg in = 1275/100 = 12.75 ml

Weight = 66 * 0.454 = 30 kg

Dose = 45 * 30 = 1350 mg

In one dose = 1350/3 = 450 mg

125 mg per ml.

450 mg in 450/125 = 3.6 ml

4 gm over 20 minutes.

So per hour = 4 * 3 = 12 gram

40 gram in 1000 ml

12 gram in 1000*12/40 = 300 ml

So rate should be 300 ml/ hour.